OpenStudy (anonymous):

A tank holds 6000 gallons of water, which drains from the bottom of the tank in half an hour. The values in the table show the volume V of water remaining in the tank (in gallons) after t minutes. t (min) 5 10 15 20 25 30 V (gal) 4116 2634 1440 606 150 0 To estimate the slope of the tangent line at t = 15, we average the slopes of the adjacent secant lines for t = 10 and t = 20. As obtained in part (a), those slopes are msec = −238.8 and msec = −166.8, respectively.

OpenStudy (anonymous):

the last part says: Therefore, the slope of the tangent line at t = 15 is as follows. (In the last step, round your answer to one decimal place.) The equation is: -238.8+(-166.8)/? =?

OpenStudy (anonymous):

part (a) had me find the Slopes of secant lines passing through P(15, 1440) and Q(t, V(t)) calculated using msec = V(t) − 1440/t − 15. It's how I found: t=10=(2634-1440/10-15)=-238.8 and t=20=(606-1440/20-15)=-166.8

OpenStudy (anonymous):

I just want to find what I have to divide by (the ?), I can do the math from there.

OpenStudy (anonymous):

it was divide by 2 and the answer was -202.8